### v1

# Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-Linear Externalities: Evidence from Triathlon using a Leave-One-Out IV Strategy

Abstract
This paper examines strategic effort and positioning choices resulting in bandwagon effects under externalities in finite multi-stage games using causal evidence from triathlon (Reichel , 2025). Focusing on open-water swim drafting—where athletes reduce drag most effectively by swimming directly behind peers—we estimate its performance effects through a structural contest framework with endogenous, deterministic effort and drafting position. Leveraging exogenous variation from COVID-19 drafting bans in Austrian triathlons, we apply a panel leave-one-out (LOO/LOTO) peer ability instrumental variables (IV) strategy to isolate the causal non-linear effect of drafting. Results from restricted sample analysis and pooled estimated bandwagon IV effects show substantial and nonlinear gains: in small (group size $<$ 10) drafting swim groups/clusters, each deeper position improves finishing rank on average by over 30%, with rapidly diminishing returns in larger groups. Leading however is consistently more costly than optimal positioning, aligning with theoretical predictions of energy expenditure (metabolic costs).
Keywords: Bandwagon effects, externalities, effort choice, contests, finite games, multi-stage games, endurance sports, triathlon, instrumental variables, drafting, swim drafting
JEL Codes: D62, L83, C26, C57
1 Introduction
Drafting plays a crucial role in race dynamics and performance outcomes in endurance and multi-endurance sports. In triathlon - where a race events consist of several finite stages, namely: swimming, cycling, and running - drafting introduces a strategic externality that greatly varies by discipline, race format and race regulation. Although drafting in professional cycling offers the most visible and studied drafting benefits, hydrodynamic drag reduction in the swim segment in triathlon are said to enhance subsequent cycling and thus overall race performance (Delextrat , 2003) offer an unique empirical setting.
This paper investigates the performance implications of open water swim drafting using rich panel data from triathlon races by modeling drafting in the first segment as a two-stage deterministic game. In the first stage, athletes exert some deterministic swimming effort; in the second stage, they position themselves optimally within a group, potentially benefiting heterogeneously from reduced drag behind other swimmers due to the inversly decaying nature of hydrodynamic benefits and rare overtakings (resorting) during the swim segment (Reichel , 2025, 2025).
Empirical estimates of drafting effects in the swim segment are still limited by measurement challenges. Group formations are likely endogenous and cannot be observed directly, and drafting exposure may correlate with unobserved ability (sorting/selection bias). Laboratory and biomechanical studies (e.g., Chatard , 1990, 2003; Delextrat , 2003; bolon2020drafting) focus on energy expenditure (metabolic) and performance gains from drafting, but large-scale causal estimates from race data have only been documented recently (in Reichel , 2025).
In our theoretical framework, we extend contest theory and incorporate models of strategic externalities (Tullock , 1980; Moldovanu , 2001; Konrad , 2009; Che , 2003) to develop a structural model in which athletes first select a deterministic level of swimming effort, followed by an endogenous choice of drafting position. Drafting benefits are captured through a non-linear function that exhibits diminishing marginal returns beyond a certain positional threshold. The structural implications of the model are then integrated into an empirical two-stage least squares (2SLS) estimation framework. To address endogeneity in within-group drafting positions, we employ a leave-one-out (LOO/LOTO) average of peer swimming ability as an instrumental variable (IV).
Our analysis provides the first large-scale heterogeneous causal estimate of open-water drafting effects using rich panel race data. This study contributes to the broader literature on strategy in multi-stage games within performance environments by offering a simple two-stage model of effort choice and externality constraints. The model aligns with frameworks that explore strategic complementarities and sequential decision-making in dynamic games (Vives , 2005; Tadelis , 2013).
2 Data
We use a rich administrative dataset provided from Triathlon Statistics Austria, covering all official triathlon races from 2010 to 2024. These include Sprint, Short, Middle, and Long distance formats, with athlete-level performance data on swim-out segment times, overall times and the final overall race ranks (after T1, cycling, T2 and the running segment).
2.1 Data Structure and Variables
The dataset is structured across three relational tables:
  • Athlete Table: Contains athlete identifiers, gender, and birth year.
  • Event Table: Includes event identifiers, calendar dates, and event category.
  • Result Table: Reports swim-out time, total race time, and final rank for each athlete-event observation.
We merge these tables using the respective foreign keys to construct a combined panel dataset indexed by athlete and event. Derived variables then include age (computed as time-dependent event year minus birth year), age squared, and binary indicators for gender and periods (pre-COVID, COVID, post-COVID). Events are grouped into pandemic periods based on their timing, capturing staggered start policies in 2020 and their subsequent relaxation into default state.
The initial dataset includes 175,740 observations. After removing records with missing values, DNFs (Did Not Finish), and DNSs (Did Not Start) athlete-level race outcomes, we obtain a cleaned panel of 168,391 athlete-event entries.
2.2 Drafting Group Construction
Drafting groups can not be observed directly in the data. We thus infer drafting clusters using a single-linkage hierarchical agglomerative clustering algorithm approach applied to swim-out segment times within each event. Athletes are grouped into the same swim cluster if their exit times fall within a 5-second distance thresholds recorded. Earlier trials with complete linkage and wider thresholds (10–15 seconds) produced overly broad or unstable groupings. (Reichel , 2025)
Within each cluster, the fastest athlete is designated as the leader, and all others are assigned ordinal positions (e.g., second drafter, third drafter) based on swim-out time segment order. We construct both continuous drafting position variables ($d_{i}$) and binary indicators (e.g., first drafter, last drafter).
2.3 Descriptive Statistics and Balance Checks
Summary statistics indicate substantial heterogeneity across events and athletes. The average swim-out time is $\approx$ 1,700 seconds, with standard deviations reflecting differences across race formats. Athletes are predominantly male ($\approx$ 79%) with a median age of $\approx$ 38. Gender and age distributions are relatively balanced across periods, though a slight aging trend in contestants is observed in the post-2020 period.
To further assess balance and validate the data quality, we compare mean swim-out times across event categories and pandemic periods (Table 1). Sprint and Short races show moderate increases in average swim-out time post-COVID, which is likely to more accurate course mapping or overall stricter timing protocols. One evidence for this is that, Middle and Long events remain more stable.
Table 1: Mean Swim-Out Times by Event Category and Pandemic Period
Category Period Mean (s) SD Observations
Sprint Pre-COVID 774 258 60,871
COVID 854 295 7,985
Post-COVID 906 296 6,152
Short Pre-COVID 1,696 415 37,722
COVID 1,855 380 4,694
Post-COVID 1,899 403 3,726
Middle Pre-COVID 2,216 436 26,711
COVID 2,311 380 4,658
Post-COVID 2,315 394 3,527
Long Pre-COVID 4,553 773 10,201
COVID 4,594 804 1,130
Post-COVID 4,572 1,035 1,014
This setup—featuring athlete-level panels, swim group formation mechanics, and quasi-random variation in drafting benefits exposure—provides a suitable empirical panel data set for testing the theoretical model introduced in subsequent Section 3.
3 Theoretical Framework
3.1 A Two-Stage Finite Game
We model drafting as a two-stage game among a finite set of athletes $\mathcal{N}=\{1,\dots,N\}$, where each athlete $i\in\mathcal{N}$ makes two sequential choices:
  1. 1.
    A swimming effort $e_{i}^{1}\in\mathbb{R}_{\geq 0}$ in the first stage;
  2. 2.
    A drafting position $d_{i}\in\mathbb{Z}_{\geq 1}$ within their drafting group in the second stage.
This structure reflects a dynamic contest framework with endogenous positioning and group-related externalities (see Tullock , 1980; Konrad , 2009; Moldovanu , 2001; Tadelis , 2013; Vives , 2005).
3.1.1 Effort Cost Function
For simplicity we assume a linear cost function of effort:
$C(e_{i}^{1})=c\cdot e_{i}^{1},\quad c>0$
3.1.2 Drafting Benefit Function
Drafting yields hydrodynamic advantages that increase with position depth beyond a threshold, reflecting growing, but the gain are diminishing up to saturation. This can be defined as following functional form:
$B(d_{i})=\begin{cases}0&\text{if }d_{i}\leq 3\\ \gamma\left(1-e^{-\lambda(d_{i}-3)}\right)&\text{if }d_{i}>3\end{cases}$
where:
  • $\gamma>0$: maximum possible drafting benefit (normalized to 1),
  • $\lambda>0$: rate of benefit growth.
Proof:
For $d_{i}>3$, it can be shown that:
$\frac{dB}{dd_{i}}>0\quad\text{(increasing in }d_{i}\text{)}$
[…]
3.1.3 Athlete’s Disutility
Athletes minimize disutility from effort (metabolic costs; likely a higher-order polynomial or exponential cost) and performance shortfalls dependent on drafting ability and/or benefits (c.p. all other externalities, hence are simply ignored in this model):
$DIS_{i}(e_{i}^{1},d_{i})=\alpha c\cdot e_{i}^{1}+(1-\alpha)(\mu-B(d_{i}))$
with:
  • $\alpha\in[0,1]$: weight on effort vs. outcome,
  • $\mu$: reference utility level (e.g., expected overall rank; a proxy such as some physiological performance predictor),
  • $B(d_{i})$: drafting benefit function.
3.2 Solving the Game: Optimal Drafting Position
Fix $e_{i}^{1}=\bar{e}_{i}^{1}$. Then the athlete’s objective becomes:
$\min_{d_{i}\in\mathbb{Z}_{\geq 1}}DIS_{i}(d_{i})=\alpha c\cdot\bar{e}_{i}^{1}-(1-\alpha)B(d_{i})$
3.2.1 Continuous Relaxation $d_{i}\in\mathbb{R}$
Assume $d_{i}\in\mathbb{R}$, $d_{i}>3$. Then:
$\frac{dDIS_{i}}{dd_{i}}<0$
Proof:
The exponential function is always positive, and $\gamma,\lambda,(1-\alpha)>0$, so the derivative is negative. Hence, disutility decreases with deeper position — but not indefinitely.
3.2.2 Second Derivative and Convexity
$\frac{d^{2}DIS_{i}}{dd_{i}^{2}}>0$
Proof:
As the second derivative is positive, disutility is strictly convex in $d_{i}$, implying a unique minimum (interior optimum).
3.2.3 Proposition (Interior Optimal Drafting Position):
Proposition 1: There exists a unique $d_{i}^{*}>3$ such that disutility is minimized, i.e., drafting benefit is maximized net of performance shortfall risk.
Proof: - For $d_{i}\leq 3$, $B(d_{i})=0\Rightarrow$ disutility is flat. - For $d_{i}>3$, $DIS_{i}$ is strictly convex. - Since $DIS_{i}\to\alpha c\bar{e}_{i}^{1}-(1-\alpha)\gamma$ as $d_{i}\to\infty$, the marginal gains vanish. - Hence, the minimum is interior and finite.
3.2.4 Empirical Mapping
We can use the assumed structural benefit function to validate its functional form, calibrate $\gamma$ and $\lambda$:
$B(d_{i})=\gamma\left(1-e^{-\lambda(d_{i}-3)}\right)\cdot\mathbf{1}\{d_{i}>3\}$
or to estimate effects. We can address address endogeneity by instrumenting $d_{i}$ using a leave-one-out peer ability IV.
3.3 Why the Optimal Position Is Not the Last
Although benefits rise in $d_{i}$, they converge to a bound:
$\lim_{d_{i}\to\infty}B(d_{i})=\gamma$
Thus, marginal benefits vanish:
$\frac{dB}{dd_{i}}=\gamma\lambda e^{-\lambda(d_{i}-3)}\to 0$
Conclusion: Athletes face trade-offs — drafting increases with depth, but so do tactical disadvantages. Therefore, the optimal position is interior — balancing maximum benefit with control. This matches observed behavior in collective movement and contests with crowding externalities (Hillman , 1989; Krueger , 1974).
4 Identification Strategy and Empirical Framework
4.1 Research Objective
We want to estimate the causal effect of within-group drafting position $D_{ieg}$ on final overall race performance $Y_{ie}$, leveraging a leave-one-out (LOO/LOTO) instrumental variable (IV) strategy. This is embedded in a finite-horizon, multi-stage game with endogenous effort, group formation. We use following notation:
  • $i$: athlete, $e$: event, $g$: drafting group;
  • $Y_{ie}$: observed performance outcome (e.g., log finishing rank + 1);
  • $D_{ieg}\in\{1,\dots,d_{\max}\}$: within-group drafting position;
  • $S_{ie}$: pre-draft swim time (a proxy for initial effort/ability);
  • $\bm{X}_{ie}$: athlete and event-level controls.
Because drafting positions reflect strategic sorting and group formation dynamics, naive OLS is likely biased. We thus use a two-stage least squares (2SLS) framework with a leave-one-out peer swim-time instrument to estimate a local average treatment effect (LATE) of drafting benefits (functional form from the theoretic model) or drafting positions on performance.
4.2 Structural Potential Outcomes Framework
Let $Y_{ie}(d)$ denote the potential outcome for athlete $i$ in event $e$ had they assumed drafting position $d$. Then:
$Y_{ie}=Y_{ie}(D_{ieg})$
Our causal estimand compares each drafting position to the leader:
$\text{ATE}(d)=\mathbb{E}[Y_{ie}(d)-Y_{ie}(1)]\quad\text{for }d>3$
Due to selection into drafting positions, unconfoundedness fails:
$\mathbb{E}[Y_{ie}\mid D_{ieg}=d]\neq\mathbb{E}[Y_{ie}(d)]$
We correct for this via an IV strategy with a theoretically derived nonlinear transformation of drafting depth.
4.3 Endogeneity and Identification Strategy
Endogeneity in drafting position arises from:
  • Strategic Sorting: Higher-ability athletes self-select into favorable positions.
  • Endogenous Grouping: Group composition reflects latent ability.
Hence:
$\text{Cov}(D_{ieg},\varepsilon_{ie})\neq 0$
We address this using a constructed leave-one-out peer swim-time instrument $Z_{ie}$ to generate the relevant and exogenous variation in swim group drafting structure.
4.3.1 Leave-One-Out (LOO/LOTO) Instrumental Variable (IV)
We, define the instrument as:
$Z_{ie}=\frac{1}{|\mathcal{G}_{ie}^{-i}|}\sum_{j\in\mathcal{G}_{ie}^{-i}}S_{je}$
This captures the swim ability of peers in athlete $i$’s drafting within the current group, excluding $i$.This leave-one-out (LOO/LOTO) instrument captures the average peer ability while avoiding mechanical reflection bias. By excluding athlete i, we remove the mechanical correlation between the instrument and the outcome. This strategy aligns with methods used in social interaction and peer effect models to address endogeneity, e.g., Bramoulle (2009) and Angrist (2014). Related applications include De Loecker (2011), who employ LOO instruments in production function estimation, and Guryan (2008), who instrument for peer characteristics using group averages excluding the individual.
4.3.2 Drafting Benefit Transformation from Theoretic Model
Instead of using raw position, we construct a structural regressor from the theory:
$B_{ieg}=\begin{cases}0&\text{if }D_{ieg}\leq 3\\ \gamma\left(1-e^{-\lambda(D_{ieg}-3)}\right)&\text{if }D_{ieg}>3\end{cases}$
We normalize $\gamma=1$, and set $\lambda=0.5$ for the sake of simplicity and brevity.
This transformation captures the increasing-but-saturating drafting benefits posited by the theoretical model.
4.4 Empirical Specification: Two-Stage Least Squares (2SLS) Peer-IV Estimation
To identify the causal effect of within-group drafting benefit on race performance, we estimate a two-stage least squares (2SLS) model using a leave-one-out peer ability instrument. This setup accounts for endogeneity in observed drafting positions, which are determined by both athlete ability and strategic group formation.
Let $i$ index athletes, $e$ index events, and $g$ index drafting clusters within an event. The key endogenous variable is $B_{ieg}$, a nonlinear transformation of the athlete’s drafting position $D_{ieg}$, derived from the structural theory. We seek to isolate exogenous variation in this drafting benefit using the average swimming ability of other athletes in the same group as an instrument.
First Stage: Instrumenting Drafting Position
We begin by estimating the drafting position as a function of peer swim ability and controls:
$D_{ieg}=\pi_{0}+\pi_{1}Z_{ie}+\pi_{2}S_{ie}+\bm{X}_{ie}^{\prime}\bm{\pi}_{3}+\lambda_{e}+\alpha_{i}+\alpha_{g}+u_{ieg}$
Where:
  • $D_{ieg}$: observed drafting position of athlete $i$ in event $e$, group $g$;
  • $Z_{ie}$: leave-one-out peer average swim time (excluding athlete $i$);
  • $S_{ie}$: own swim time of athlete $i$, capturing own ability;
  • $\bm{X}_{ie}$: vector of additional athlete- or event-level covariates (e.g., age, gender);
  • $\lambda_{e}$: event fixed effects, capturing course characteristics and environmental conditions;
  • $\alpha_{i}$: athlete fixed effects, absorbing time-invariant ability heterogeneity;
  • $\alpha_{g}$: group fixed effects, accounting for shared cluster-level dynamics;
  • $u_{ieg}$: idiosyncratic first-stage error term.
The instrument $Z_{ie}$ introduces exogenous variation by exploiting group composition. Since athletes do not control who else is in their drafting cluster, this measure captures plausibly exogenous shifts in strategic positioning pressure.
Second Stage: Estimating the Effect of Drafting Benefit on Performance
Using the fitted drafting benefit $\widehat{B}_{ieg}=f(\widehat{D}_{ieg})$ to validate the theoretical model or using $\widehat{D}_{ieg})$ directly, we estimate:
$\log(rank_{ie}+1)=\beta_{1}\widehat{B}_{ieg}+\beta_{2}\cdot leader_{ieg}+\bm{X}_{ie}^{\prime}\bm{\beta}_{3}+\lambda_{e}+\alpha_{i}+\alpha_{g}+\varepsilon_{ieg}$
Where:
  • $\log(rank_{ie}+1)$: log-transformed within-event finishing rank of athlete $i$, adjusted to avoid log-zero;
  • $\widehat{B}_{ieg}$: instrumented drafting benefit, based on the theoretical transformation of $\widehat{D}_{ieg}$;
  • $leader_{ieg}$: indicator variable for being in the frontmost drafting position within group $g$;
  • $\bm{X}_{ie}$, $\lambda_{e}$, $\alpha_{i}$, $\alpha_{g}$: as defined above;
  • $\varepsilon_{ieg}$: second-stage error term.
Identification Strategy and Exogeneity
This 2SLS framework identifies the local average treatment effect (LATE) of drafting benefit for compliers—athletes whose position is affected by peer group composition. Identifying Assumptions are:
  • Instrument relevance: $\text{Cov}(Z_{ie},D_{ieg})\neq 0$;
  • Instrument exogeneity: $\mathbb{E}[\varepsilon_{ieg}\mid Z_{ie},S_{ie},\bm{X}_{ie},\lambda_{e},\alpha_{i},\alpha_{g}]=0$. – $Z_{ie}$;
  • Monotonicity: An increase in peer ability (faster peers) weakly shifts athletes to deeper drafting positions.
This empirical structure provides consistent estimates of the causal effect of drafting benefit on race outcomes while accounting for sorting, selection, and unobserved heterogeneity through fixed effects and instrument design.
where: - $\widehat{B}_{ieg}$: instrumented drafting benefit; - $leader_{ieg}$: indicator for leading position ($D=1$); - $\lambda_{e},\alpha_{i},\alpha_{g}$: event, athlete, and group fixed effects.
4.4.1 Structural Validation and Interpretation
The coefficient $\beta_{1}$ reflects the average marginal return to drafting benefit. The increasing and concave structure implies that: - Greatest positive net gains occur when transitioning from front positions backwards; Diminishing returns occur deeper into the pack; The optimal position is interior, as predicted by the theoretical model.
Empirical evidence of this shape supports the game’s interior equilibrium logic and nonlinear (dis-)utility (drafting befenefit structure) topology.
4.5 Potential Extension: Projected Panel IV
For robustness and within-athlete variation, we define a projected instrument using historical group characteristics from the largest group containing athlete ’i’ f.e.:
$Z_{ie^{\prime}}^{\text{histProjected}}=\mathbb{E}_{e}[Z_{ie}\mid i\in g_{ie},e\neq e^{\prime}]$
This makes more use of our rich panel data set with repeated athletes across multiple events to increase within-athlete variation.
5 Empirical Results
Restricting the analysis to the capped/limited outcome specification, where the dependent variable is transformed as $\log(r_{i}+1)$ and $r_{i}<250$, we isolate performance effects among most finishers (compliers) and obtain cleaner identification (robust to overly large outliers) of drafting benefit dynamics. The instrumental variables (IV) estimates consistently show that deeper drafting positions improve performance, with stronger estimated effects in smaller groups.
In Table 3, the coefficient on drafting position is $-0.011$ ($p<0.1$) in the full sample (without restricting the outcome variable), implying that moving one position deeper in the drafting line improves expected rank by approximately $1.1\%$:
%$\%\Delta r_{i}\approx(e^{-0.011}-1)\times 100\approx-1.1\%$
This effect increases substantially in smaller groups. In drafting clusters with fewer than 10 athletes, the coefficient becomes $-0.972$ ($p<0.01$), translating to a $62.1\%$ improvement in finishing rank per position:
%$\%\Delta r_{i}\approx(e^{-0.972}-1)\times 100\approx-62.1\%$
In very small clusters (group size $<5$), the coefficient is $-2.378$ ($p<0.1$), corresponding to an $90.7\%$ improvement:
%$\%\Delta r_{i}\approx(e^{-2.378}-1)\times 100\approx-90.7\%$
These patterns align closely with the theoretical model, which states steep but diminishing marginal benefits from drafting due to overall saturation effects.
5.1 Using the theory-based functional form
The theory-based specifications in Table 6 reinforce these findings using an improved drafting benefit function:
$B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ 1-e^{-\lambda(d_{i}-3)},&\text{if }d_{i}>3\end{cases}$
with some uncalibrated $\lambda=0.5$ for the simple sake of simplicity and brevity. This specification reflects increasing but diminishing returns to deeper drafting positions. Across all models, the coefficient on $B(d_{i})$ is significantly negative, consistent with improved performance from stronger drafting effects. The effect is strongest in smaller groups: in groups with fewer than 10 athletes, a one-unit increase in drafting benefit improves log-rank by approximately 7.1 to 9.6 points (Columns 5–6). In broader samples, the benefit remains significant and negative—$-6.65$ in the full sample (Column 1), and $-4.84$ to $-3.61$ in groups smaller than 20 (Columns 3–4).
Estimates on the Leader dummy are also consistently negative and statistically significant where included, ranging from $-1.65$ to $-2.85$. These results suggest that athletes at the front of their drafting group perform better on average. Given the IV framework controls for endogenous group sorting, the leader coefficient likely captures small/negligible non-effort-based advantages such as tactical control, friction or reduced group congestion—complementing prior evidence on physiological pacing costs (Menting , 2022).
All models reject exogeneity via the Wu–Hausman test ($p<2.2\times 10^{-16}$), validating the use of instrumental variables. First-stage F-statistics confirm strong instrument relevance across all samples, with particularly robust strength in the full sample ($F=254.7$), the alternate IV specification ($F=513.7$), and small-group settings ($F=99.1$ to $148.5$).
6 Conclusion
This paper contributes comprehensive analysis on Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-linear Externalities and draws on empirical causal evidence using rich panel data from drafting in endurance sports competitions, predicting effects on final stage rank outcomes focusing on competitive athletes. We develop and estimate log-linear models with endogenous drafting position, using a leave-one-out peer swim ability as an instrument (Angrist , 2009; Manski , 1993; De Giorgi , 2020). The results validate the strategic and nonlinear nature of drafting benefits in multi-stage competitions such as triathlon.
Our key findings are:
  1. 1.
    Drafting deeper in the pack improves performance: In the full sample, a one-position drop in the drafting line improves expected finishing rank by $1.1\%$ ($\hat{\beta}=-0.011$). In small groups (size $<10$), this effect grows to $62.1\%$ ($\hat{\beta}=-0.972$, $p<0.01$). In very small groups ($<5$), the effect reaches $90.7\%$ ($\hat{\beta}=-2.378$, $p<0.1$).
  2. 2.
    Theory-based drafting benefits follow nonlinear decay: In smaller groups, a one-unit increase in uncalibrated theoretical drafting benefit $B(d_{i})$ improves performance by 22–30 log-rank points. These results validate nonlinear decay in marginal gains beyond the third drafting position, in line with the optimal interior solution from the theoretic model.
  3. 3.
    Leading is costly: While leader coefficients are negative, IV estimates in smaller clusters propose that leading imposes performance costs due to higher energy expenditure—reinforcing the value of tactical positioning behind the leader (Delextrat , 2003).
  4. 4.
    Pooled bandwagon effects highlight the stylized fact that overall race performance improves with latter drafting position: In a series of binary IV comparisons across adjacent drafting position bands (e.g., 1–2 vs 3–4, 2–3 vs 4–5), we find large, statistically significant reductions in adjusted rank—ranging from 31.8% to 60.3%—associated with moving into a latter drafting position group. That is, athletes further back in the pack tend to finish with better ranks, consistent with drafting-based performance spillovers. However, this benefit plateaus at some point and becomes statistically insignificant beyond positions 7–8, indicating a capped or diminishing bandwagon effect or aligned with the in theory derived interior optimum.
By focusing on capped/limited outcome ranks, we reduce the influence of extreme values and measurement error in the fat tails. The findings offer strong empirical support for contest models with externalities (Tullock , 1980; Che , 2003) and provide new insights into strategy-performance interactions in spatially structured environments. These results have implications not only for endurance sports but also for organizational settings where individuals compete in the presence of similar externalities and strategic effort in multi-stage games with finite horizon.
Appendix A Appendix
Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-Linear Externalities: Evidence from Triathlon using a Leave-One-Out IV Strategy
A.1 Leave-one-out Peer-IV 2SLS Regressions with Capped/Limited Outcome Variable Rank $<$ 250
Table 2: Instrumental Variables Estimates of Drafting Effects on Performance
(1) Full Sample (2) Group Size $<$ 15 (3) Group Size $<$ 10 Drafting Position (Fit) -0.039∗∗∗ -1.729∗∗∗ -2.496∗∗ (0.007) (0.491) (0.910) Leader -0.301∗∗∗ -5.328∗∗∗ -6.159∗∗ (0.052) (1.526) (2.235) Observations 45,215 26,357 21,190 Athletes (FE) 15,762 11,919 10,534 Events (FE) 779 793 786 Clusters (FE) 149 149 149 RMSE 1.036 3.601 3.339 Adj. $R^{2}$0.128 -11.8 -13.6 Within $R^{2}$1.979 -33.1 -46.8 First Stage F-stat 998.4 24.7 14.6 Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Estimates are based on two-stage least squares (2SLS), with drafter position treated as endogenous and instrumented by the leave-one-out cluster-level mean swimming ability. The dependent variable is $\log(r_{i}+1)$, where $r_{i}$ denotes the centered within-event finishing rank of athlete $i$. All models include fixed effects for athlete, event, and drafting cluster. Standard errors are clustered at the event level. In Column (1), the coefficient on drafter position is small but significantly negative, suggesting that being farther behind in a drafting cluster is associated with better performance (lower log rank). The effect increases substantially in magnitude and significance in Columns (2) and (3), which restrict the sample to smaller clusters, implying stronger drafting dynamics in such settings. The coefficient on the Leader dummy is also consistently negative and statistically significant, indicating that athletes who lead clusters tend to finish with better ranks. This likely reflects a combination of superior ability and favorable positioning strategy. A total of 78,384 (full sample), 50,973 (group size $<$ 15), and 42,541 (group size $<$ 10) observations were excluded due to missing or invalid values of the dependent variable, including cases where $\log(r_{i}+1)$ was undefined or produced NaNs. This may stem from negative centering; future versions could offset $r_{i}$ for robustness. Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
Table 3: Instrumental Variables Estimates of Interpretable Drafting Effects on Performance
(1) Full Sample (2) Group Size $<$ 10 (3) Group Size $<$ 5 Drafter Position (Fit) -0.011 -0.972∗∗∗ -2.378 (0.0053) (0.185) (1.073) Observations 34,061 16,035 9,316 Athletes (FE) 12,181 8,408 5,856 Events (FE) 438 523 531 Clusters (FE) 178 178 176 RMSE 0.771 1.161 1.176 Adj. $R^{2}$ 0.641 -0.166 -0.674 Within $R^{2}$0.905 -5.925 -8.453 First Stage F-stat 1,115.6 51.4 12.7 First Stage $p$-value $<2.2\times 10^{-16}$ $8.23\times 10^{-13}$ $3.64\times 10^{-4}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the centered finishing rank of athlete $i$ within event. The endogenous regressor is the athlete’s drafter position $p_{i}$, instrumented using the cluster-level peer mean swim performance (leave-one-out). No additional covariates are included. Fixed effects for athlete, cluster, and event are used. Standard errors are clustered at the event level. In Column (1), the coefficient is small but statistically significant, indicating modest benefits from trailing farther behind. The magnitude and significance of the effect increase in Columns (2) and (3), which restrict to smaller drafting clusters. In these subsamples, being positioned farther back in the pack is associated with significantly better performance, consistent with drafting theory. The precision of the estimates improves in Column (2) due to stronger first-stage relevance (F = 51.4), though instrument strength is more limited in Column (3) (F = 12.7). All models pass the Wu-Hausman test, supporting the endogeneity of drafter position. A total of 52,222 and 30,113 observations were dropped in Columns (2) and (3), respectively, due to missing or invalid values of the dependent variable, including cases where $\log(r_{i}+1)$ is undefined (e.g., due to negative centering). Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
A.2 Robustness: Leave-one-out Peer-IV 2SLS Regressions under different sample sizes and model specifications
Table 4: The Causal Effects of Drafting Position on Log Race Rank (2SLS Estimates)
Capped Drafting Position Uncapped Drafting Position (1) Full FE (2) +Athlete FE (3) +Event FE (4) +Cluster FE (5) +Athlete FE Drafting Position (Fit) $-0.504^{***}$ $-4.741^{*}$ $-0.046^{***}$ $-0.083^{***}$ $-0.011^{*}$ (0.069) (2.321) (0.011) (0.013) (0.005) Leader Indicator $-1.889^{***}$ $-6.425^{*}$ $-0.797^{***}$ $-0.848^{***}$ $-0.103^{*}$ (0.261) (3.159) (0.117) (0.137) (0.048) Observations 23,544 9,316 34,061 34,061 34,061 RMSE 1.34 1.42 3.20 5.22 0.78 Adjusted $R^{2}$0.30 -1.43 -2.90 -9.46 0.64 Within $R^{2}$6.06 -12.70 -8.52 -24.80 -0.94 First-stage $F$ 87.3 8.42 400.3 129.6 1,076.5 Wu–Hausman 530.9 93.0 3,755.8 3,518.6 1,038.0 Athlete Fixed Effects Event Fixed Effects Cluster Fixed Effects Notes: All models are estimated via two-stage least squares (2SLS), instrumenting Drafting Position using the leave-one-out mean swim ability of other swimmers in the drafting group. The dependent variable is $\log(\text{rank}+1)$, a log-transformation of race placement. Columns (1)–(2) use a version of Drafting Position capped at 5 to mitigate outlier influence; Columns (3)–(5) use the uncapped measure. Event fixed effects are included in all models. Athlete fixed effects appear in Columns (2) and (5); cluster fixed effects in Columns (4) and (5). Standard errors are clustered at the event level. Interpretation: Coefficients represent semi-elasticities: a one-position increase in drafting position (i.e., being farther behind) changes expected finishing rank by approximately $100\cdot[\exp(\beta)-1]\%$: (1) $\beta=-0.504\Rightarrow$  39.7% increase in rank (Strong Instrument, Rank Capped/Outliers $>$ 250 removed) (2) $\beta=-4.741\Rightarrow$  374% increase (Weak Instrument!) (5) $\beta=-0.011\Rightarrow$  1.1% increase with full controls (Fixed Effects) The coefficient on Leader Indicator reflects the cost of leading early. Negative estimates suggest that being the lead swimmer may reduce final performance, possibly due to pacing fatigue or exposure. First-stage $F$-statistics exceed the conventional threshold of 10 in all but Column (2), confirming instrument strength. Wu–Hausman tests reject exogeneity of Drafting Position, supporting a causal interpretation. Under standard IV assumptions (relevance, exclusion, monotonicity), the estimates identify the Local Average Treatment Effect (LATE) of drafting position on log-transformed race outcomes. Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
A.3 Robustness: Leave-one-out Peer-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function
Table 5: Instrumental Variables Estimates of Theory-Based Drafting Effects on Performance
(1) Group Size $>$ 1 (2) Group Size $<$ 20 (3) Group Size $<$ 10 Drafting Benefit (Fit) -6.650∗∗∗ -10.662∗∗∗ -16.711∗∗ (0.654) (1.871) (5.239) Leader -2.845∗∗∗ -3.676∗∗∗ -3.980∗∗ (0.288) (0.659) (1.256) Observations 45,215 29,073 21,190 Athletes (FE) 15,762 12,544 10,534 Events (FE) 779 789 786 Clusters (FE) 149 149 149 RMSE 1.717 2.570 3.165 Adj. $R^{2}$2.100 -6.635 -12.100 Within $R^{2}$7.186 -22.200 -42.000 First Stage F-stat 254.7 53.0 16.3 First Stage $p$-value $<2.2\times 10^{-16}$ $3.54\times 10^{-13}$ $5.41\times 10^{-5}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Each column reports two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the athlete’s centered within-event rank. The endogenous variable is a theory-based drafting benefit: $B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ \gamma\left(1-e^{-\lambda(d_{i}-3)}\right),&\text{if }d_{i}>3\end{cases}$ with $\gamma=1$ and $\lambda=0.5$. The instrument is the cluster-level peer mean swim time (leave-one-out). Fixed effects for athlete, cluster, and event are included. Standard errors are clustered at the event level. In Column (1), covering all valid group sizes, the drafting benefit significantly reduces disutility (better performance), with a very strong first-stage instrument. Effects are slightly larger in magnitude in Columns (2) and (3), reflecting increased sorting and clearer benefit signals in smaller, more cohesive drafting groups. The leader dummy is negative and significant in all samples, suggesting that leading is correlated with stronger athletes, consistent with selection or effort asymmetries. Log-rank values may include centering; NA values and log issues led to removal of 78,384; 56,239; and 42,541 observations in Columns (1) to (3) respectively. Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
Table 6: 2SLS Estimates Using Theory-Implied Drafting Benefit $B(d_{i})=1-e^{-\lambda(d_{i}-3)}$
(1) (2) (3) (4) (5) (6) Group $>$ 1 Group > 1, No Leader Group $<$ 20 Group $<$ 20, No Leader Group $<$ 10 Group $<$ 10, No Leader Theoretic Drafting Benefit Fit $B(d_{i})$ $-6.650^{***}$ $-2.871^{***}$ $-4.836^{***}$ $-3.609^{***}$ $-9.615^{***}$ $-7.069^{***}$ (0.654) (0.267) (0.697) (0.468) (2.371) (1.509) Leader Indicator $-2.845^{***}$ $-1.645^{***}$ $-2.053^{***}$ (0.288) (0.242) (0.511) Observations 45,215 34,061 22,026 22,026 16,035 16,035 Athletes (FE) 15,762 12,181 10,075 10,075 8,408 8,408 Events (FE) 779 438 498 498 523 523 Clusters (FE) 149 178 178 178 178 178 RMSE 1.717 0.925 1.201 1.020 1.661 1.339 Adj. $R^{2}$2.10 0.483 -0.075 0.225 -1.383 -0.550 Within $R^{2}$7.19 -1.75 -5.13 -3.42 -13.20 -8.21 First-stage $F$-stat 254.7 513.7 99.1 148.5 23.4 37.4 Wu–Hausman stat 1,901.0 918.2 509.1 508.6 298.5 297.7 Athlete FE Event FE Cluster FE Notes: 2SLS estimates of $\log(r_{i}+1)$, where $r_{i}$ is centered rank. The endogenous regressor is the theory-based drafting benefit: $B(d_{i})=1-e^{-\lambda(d_{i}-3)}$ for $d_{i}>3$, and 0 otherwise, with $\lambda=0.5$. Instrument: leave-one-out peer swim time. Fixed effects for athlete, event, and cluster are included. Standard errors clustered at the event level. Leader coefficients are shown where included. Column labels correspond to different subsamples and whether the leader dummy is included. Drafting benefits are consistently significant and negative, with stronger effects in smaller groups. Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$
Appendix B Data and Descriptive Statistics as in (Reichel , 2025)
Variable Min 1st Quartile Median Mean 3rd Quartile Max Total 165 4965 8232 11789 16809 61282 Swim-Out Times 4 825 1431 1625 2097 29700 Race Rank 0 28 69 156.9 151 2623 Male 0.0000 1.0000 1.0000 0.7898 1.0000 1.0000 Year 1922 1969 1977 1977 1985 2009 Covid 0.0000 0.0000 0.0000 0.1097 0.0000 1.0000 Post Covid 0.00000 0.00000 0.00000 0.08563 0.00000 1.00000 Week Running 0.0 159.0 279.0 317.6 475.0 751.0 Age 8.00 31.00 38.00 38.56 46.00 99.00 Age Squared 64 961 1444 1600 2116 9801 Event Year 2010 2013 2015 2016 2019 2024 Cluster (Swim Group) 1.00 4.00 10.00 16.95 23.00 224.00 Leader 0.0000 0.0000 0.0000 0.3134 1.0000 1.0000 Drafter 0.0000 0.0000 1.0000 0.6866 1.0000 1.0000 Drafter Position 1.00 1.00 3.00 17.15 11.00 937.00 First Drafter 0.0000 0.0000 0.0000 0.2961 1.0000 1.0000 Second Drafter 0.0000 0.0000 0.0000 0.1342 0.0000 1.0000 Third Drafter 0.00000 0.00000 0.00000 0.08329 0.00000 1.00000 Fourth Drafter 0.00000 0.00000 0.00000 0.05846 0.00000 1.00000 Fifth Drafter 0.000 0.000 0.000 0.044 0.000 1.000 Last Drafter 0.0000 0.0000 0.0000 0.2961 1.0000 1.0000 Notes: This table presents summary statistics for the key variables used in the analysis. The dataset consists of 168,391 observations. Variables include athlete characteristics, race details, and encoded swim group (cluster), drafting, drafting position binary dummy variables. Encoded categorical variables such as Event Category $\in\{\text{Sprint},\text{Short},\text{Middle},\text{Long}\}$ are not displayed.
Table 7: Summary Statistics of numerical variables
Table 8: Balance Check of Swim-Out Times across different Event Categories and Periods
Event Category Period Mean Swim-Out Time SD Swim-Out Time Count
Short Covid 1855 380 4694
Short Post-Covid 1899 403 3726
Short Pre-Covid 1696 415 37722
Long Covid 4594 804 1130
Long Post-Covid 4572 1035 1014
Long Pre-Covid 4553 773 10201
Middle Covid 2311 380 4658
Middle Post-Covid 2315 394 3527
Middle Pre-Covid 2216 436 26711
Sprint Covid 854 295 7985
Sprint Post-Covid 906 296 6152
Sprint Pre-Covid 774 258 60871
Table 9: Balance Check across periods
Period Mean Total Mean Swimming Time Mean Rank Mean Age Mean Male Count Covid 12091 1705 121 40.2 0.768 18467 Post-Covid 12573 1765 127 40.8 0.757 14419 Pre-Covid 11665 1599 165 38.1 0.796 135505
Figure 1: Age Distribution by Period Balance Check
Figure 2: Swim Out Times by Event Category and Period Balance Check
Appendix C Robustness: OLS Estimations using Within-Group Drafter Position Dummies (as in (Reichel , 2025))
Table 10: OLS Regression Results Considering all Participants in Non-Leading Swim-Group Roles as Potential Drafting Beneficiaries
(1) Pre-Covid Sample ($\leq 2020$) (2) Covid Sample (2020, 2021, 2022) (3) Post-Covid Sample ($\geq 2023$) (4) Full Sample (2010 - 2024)
Panel A: OLS Estimations (Swim Group Def.: Max 5 sec apart, Cluster (Swim Group) Complete Linkage, Subsample based on Athletes: Competing in all 3 Periods)
Observations 29,750 9,796 6,941 46,487
Fixed Effects Event, Athlete Event, Athlete Event, Athlete Event, Athlete
Clustered SE Event ID Event ID Event ID Event ID
Drafter -19.476*** -13.061*** -15.648** -18.569***
Standard Error (2.413) (3.517) (4.668) (1.855)
RMSE 171.3 147.9 129.7 178.9
Adjusted R2 0.9709 0.9742 0.9793 0.9698
Within R2 0.0021 0.0012 0.0020 0.0019
Panel B: OLS Estimations (Swim Group Def.: Max 5 sec apart, Single Swimmers are allowed, Cluster (Swim Group) Complete Linkage)
Observations 135,505 18,467 14,419 168,391
Fixed Effects Event, Athlete Event, Athlete Event, Athlete Event, Athlete
Clustered SE Event ID Event ID Event ID Event ID
Drafter -22.892*** -18.307*** -22.388*** -23.047***
Standard Error (1.459) (4.179) (4.365) (1.330)
RMSE 166.6 140.1 116.1 179.4
Adjusted R2 0.9717 0.9707 0.9781 0.9668
Within R2 0.0035 0.0024 0.0045 0.0030
Panel C: OLS Estimations (Swim Group Def.: Max 15 sec apart, Cluster (Swim Group) Complete Linkage, Subsample based on Athletes: At Least 3 Swimmers per Group)
Observations 111,880 14,262 11,011 137,153
Fixed Effects Event, Athlete Event, Athlete Event, Athlete Event, Athlete
Clustered SE Event ID Event ID Event ID Event ID
Drafter -5.348*** -3.206 -9.764* -5.339***
Standard Error (0.951) (2.620) (4.082) (0.841)
RMSE 122.7 92.1 76.5 123.4
Adjusted R2 0.9809 0.9826 0.9860 0.9803
Within R2 0.00024 0.00013 0.00139 0.00024
Significance Codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$
Panel D: OLS Estimation (Dependent Variable: Swim-Out Time, Clustered SE at Athlete Level, Swim Group Def.: Max 5 sec. apart, Cluster (Swim Group) Single Linkage)
Observations 168,391
Fixed Effects Athlete (29,194), Event (1,339)
Clustered SE Athlete ID
Variable Estimate Std. Error t value Pr(>|t|)
Pre-period x Drafter -51.680*** 1.841 -28.065 < 2.2e-16
Post-period x Drafter 15.138*** 3.629 4.171 3.04e-05
RMSE 178.8
Adjusted R2 0.9670
Within R2 0.0099
Panel E: OLS Estimation (Dependent Variable: Swim-Out Time, Additional Controls, Swim Group Def.: Max 5 sec. apart, Cluster (Swim Group) Single Linkage)
Observations 168,391
Fixed Effects Athlete (29,194), Event (1,339)
Clustered SE Athlete ID
Pre-period x Drafter -29.858*** 1.880 -15.886 < 2.2e-16
Post-period x Drafter 14.012*** 3.627 3.864 0.00011
Drafter Position -0.173*** 0.018 -9.398 < 2.2e-16
Cluster (Swim Group) 1.669*** 0.107 15.564 < 2.2e-16
RMSE 177.4
Adjusted R2 0.9675
Within R2 0.0251
Panel F: OLS Estimation (Dependent Variable: Swim-Out Time, Clustered SE at Athlete & Event Level, Swim Group Def.: Max 5 sec. apart, Cluster (Swim Group) Single Linkage)
Observations 168,391
Fixed Effects Athlete (29,194), Event (1,339)
Clustered SE TWFE (Athlete ID, Event ID)
Pre-period x Drafter -29.967*** 2.433 -12.317 < 2.2e-16
Post-period x Drafter 12.543** 4.357 2.879 0.0041
Drafter Position -0.271** 0.091 -2.978 0.0030
Cluster (Swim Group) 1.518*** 0.124 12.285 < 2.2e-16
Race Rank 0.402*** 0.030 13.316 < 2.2e-16
RMSE 169.4
Adjusted R2 0.9704
Within R2 0.1111
Significance codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$
Panel G: OLS Estimation (Additional Swim Group Cluster-Level Fixed Effects, Robust SE, Swim Group Def.: Max 5 sec. apart, Cluster (Swim Group) Single Linkage)
Fixed Effects Cluster (Swim Group) (224), Athlete (29,194), Event (1,339)
Clustered SE Robust SE (HC1)
Pre-period x Drafter -31.401*** 1.652 -19.013 < 2.2e-16
Post-period x Drafter 12.562*** 3.594 3.495 0.00047
Drafter Position -0.180*** 0.018 -10.277 < 2.2e-16
RMSE 175.0
Adjusted R2 0.9684
Within R2 0.0045
Significance codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$
Table 11: OLS Estimation Results for Swim-Out Times Under Different Model Specifications
Notes: This table presents results from OLS regressions estimating the effects of drafting on Swim-out times performance. (lower is better) Panel D estimates the effects using an interaction terms specification, clustering at the athlete level. Panel E introduces additional controls such as the swimmer’s relative position in the drafting group ($drafter\_position$) and cluster ID denoting the swim group. Panel F further refines the estimation by clustering standard errors at both the athlete and event level. The dependent variable is the Swim-Out Time time in all models. Panel G extends the analysis by including cluster-level fixed effects, which account for within-group variations in drafting performance, while using HC1 standard errors robust to heteroskedasticity.
Significance codes: *** $p\leq 0.001$, ** $p\leq 0.01$, * $p\leq 0.05$.
Table 12: Drafting Position Dummies FE Model Estimation Results
Estimate Std. Error t value Pr(>|t|)
First Drafter 71.9698 1.72555 41.70837 < 2.2e-16 ***
Second Drafter 39.4901 1.83705 21.49654 < 2.2e-16 ***
Third Drafter 28.1694 1.84785 15.24444 < 2.2e-16 ***
Fourth Drafter 23.5146 2.02665 11.60269 < 2.2e-16 ***
Fifth Drafter 19.5379 2.08455 9.37268 < 2.2e-16 ***
  • Notes: This table presents the results of an ordinary least squares (OLS) estimation with fixed effects. The dependent variable is Swim-Out Time (sec). The model includes fixed effects for athlete_id (29,194 levels) and event_id (1,339 levels). Robust standard errors are used to account for heteroskedasticity.
  • The number of observations is 168,391. The root mean squared error (RMSE) is 178.5. The adjusted $R^{2}$ is 0.967166, and the within-group $R^{2}$ is 0.013724.
  • Statistical significance codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Table 13: Contd. Drafting Position Dummies FE Model Estimation Results
Estimate Std. Error t value Pr(>|t|)
Panel A: OLS Estimation FE Model with Drafting Positions
First Drafter 71.9698 1.72555 41.70837 < 2.2e-16 ***
Second Drafter 39.4901 1.83705 21.49654 < 2.2e-16 ***
Third Drafter 28.1694 1.84785 15.24444 < 2.2e-16 ***
Fourth Drafter 23.5146 2.02665 11.60269 < 2.2e-16 ***
Fifth Drafter 19.5379 2.08455 9.37268 < 2.2e-16 ***
Panel B: OLS Estimation FE Model with Drafting Positions and Additional Controls
Leader -26.7843 4.35217 -6.15423 7.5649e-10 ***
Cluster (Swim Group) 0.65793 0.08622 7.63061 2.3514e-14 ***
Race Rank 0.39668 0.00723 54.85970 < 2.2e-16 ***
First Drafter 48.5491 4.23790 11.45593 < 2.2e-16 ***
Second Drafter 31.2118 1.79029 17.43395 < 2.2e-16 ***
Third Drafter 23.5610 1.74702 13.48643 < 2.2e-16 ***
Fourth Drafter 19.6363 1.89143 10.38174 < 2.2e-16 ***
Fifth Drafter 15.4568 1.93864 7.97301 1.5604e-15 ***
Leader x Cluster (Swim Group) 1.29262 0.12643 10.22384 < 2.2e-16 ***
  • Notes: This table presents the results of two ordinary least squares (OLS) estimations with fixed effects. The dependent variable is Swim-Out Times. The models include fixed effects for athlete_id (29,194 levels) and event_id (1,339 levels). Robust standard errors are used to account for heteroskedasticity.
  • The number of observations is 168,391. Model 1 has a root mean squared error (RMSE) of 178.5, an adjusted $R^{2}$ of 0.967166, and a within-group $R^{2}$ of 0.013724. Model 2 has an RMSE of 169.2, an adjusted $R^{2}$ of 0.970474, and a within-group $R^{2}$ of 0.113118.
  • Statistical significance codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Table 14: Combined Drafting Position Dummies FE Model Estimation Results
Model 1 Model 2 Model 3
Leader - -26.7843∗∗∗ 0.8688
(4.3522) (3.9616)
Cluster (Swim Group) - 0.6579∗∗∗ -
(0.0862) -
Race Rank - 0.3967∗∗∗ -
(0.0072) -
First Drafter 71.9698∗∗∗ 48.5491∗∗∗ 46.8571∗∗∗
(1.7256) (4.2379) (4.3806)
Second Drafter 39.4901∗∗∗ 31.2118∗∗∗ 26.0330∗∗∗
(1.8371) (1.7903) (1.8769)
Third Drafter 28.1694∗∗∗ 23.5610∗∗∗ 19.4259∗∗∗
(1.8479) (1.7470) (1.8238)
Fourth Drafter 23.5146∗∗∗ 19.6363∗∗∗ 17.4769∗∗∗
(2.0267) (1.8914) (1.9906)
Fifth Drafter 19.5379∗∗∗ 15.4568∗∗∗ 15.3214∗∗∗
(2.0846) (1.9386) (2.0454)
Observations 168,391 168,391 168,391
RMSE 178.5 169.2 175.0
Adj. $R^{2}$ 0.9672 0.9705 0.9684
Within $R^{2}$ 0.0137 0.1131 0.0050
  • Notes: This table presents the results of three fixed-effects OLS estimations. The dependent variable is Swim-Out Times. Models include fixed effects for athlete_id (29,194 levels) and event_id (1,339 levels). Standard errors are robust to heteroskedasticity.
  • Statistical significance: *** $p<0.001$, ** $p<0.01$, * $p<0.05$.
Table 15: OLS Estimation Results - Dependent Variable: Swim Out Times
Estimate Std. Error t-value Pr(>|t|)
Drafter -23.8886 4.1184 -5.8004 $6.63\times 10^{-9}$**
Pre-period $\times$ Drafter -8.5779 4.6013 -1.8642 0.0623 .
Covid-year $\times$ Drafter 19.4150 8.1469 2.3831 0.0172 *
Observations 168,391
Fixed-effects Athlete ID: 29,194, Event ID: 1,339, Cluster (Swim Group): 224
Standard-errors Heteroskedasticity-robust
RMSE 175.1
Adj. $R^{2}$ 0.9683
Within $R^{2}$ 0.0035
  • Notes:
  • The dependent variable is Swim-Out Times performance (time in seconds). Fixed-effects regression includes athlete ID, event ID, and computed swim group cluster in event.
  • Periods: The variable pre-period is a dummy indicating events before 2020. The variable covid-year is a dummy indicating events in 2020.
  • Swim groups: Clusters are determined by hierarchical clustering with a threshold of 5 seconds.
  • Dummies: The drafter dummy is 1 for non-leaders of swim groups. The pre-period * drafter interaction tests the effect of drafting before 2020. The covid-year * drafter interaction examines the drafting effect in 2020.
  • Significance codes: *** $p<0.001$, ** $p<0.01$, * $p<0.05$, . $p<0.1$.
Appendix D Pooled Bandwagon Effects of Latter Drafting Positions on Overall Rank
To estimate the causal effect of drafting position on relative athlete performance, we specify a two-stage least squares (2SLS) model with a log-linear functional form. The structural equation of interest is:
$\log(Rank_{it}^{*})=\alpha+\beta\cdot\text{Treat}_{it}+\gamma_{i}+\delta_{t}+\varepsilon_{it}$ (1)
where:
  • $Rank_{it}^{*}=Rank_{it}-\bar{Rank}_{t}+c$ is the event-centered rank for athlete $i$ in event $t$, shifted by a constant $c>0$ to ensure positivity.
  • $\text{Treat}_{it}\in\{0,1\}$ is a binary variable equal to 1 if the athlete is in the worse of two pooled drafting positions (e.g., positions 3–4 rather than 1–2).
  • $\gamma_{i}$ and $\delta_{t}$ are cluster and event fixed effects, respectively.
  • $\varepsilon_{it}$ captures unobserved performance shocks.
Because $\text{Treat}_{it}$ may be endogenous—potentially correlated with $\varepsilon_{it}$—we instrument for it using the exogenous variation in peer group swim ability. The first-stage equation is:
$\text{Treat}_{it}=\pi\cdot\text{PeerSwimMean}_{it}+\eta_{i}+\theta_{t}+u_{it}$ (2)
where:
  • $\text{PeerSwimMean}_{it}$ is the leave-one-out average swim score of athlete $i$’s drafting cluster in event $t$.
  • $\eta_{i}$, $\theta_{t}$ are the same fixed effects, and $u_{it}$ is the first-stage error term.
Under the standard IV assumptions—instrument relevance ($\pi\neq 0$), instrument exogeneity ($\mathbb{E}[\text{PeerSwimMean}_{it}\cdot\varepsilon_{it}]=0$), and monotonicity—the parameter $\beta$ captures the local average treatment effect (IV-LATE) of being a drafter in latter pooled drafter position band on log-transformed, adjusted rank.
Results
Table 16: Instrumental Variable Estimates: Effect of Drafter Position on Adjusted Rank
Comparison Estimate (log pts) Std. Error Lower CI Upper CI P-value % Change
1–2 vs 3–4 $-0.382$ $0.031$ $-0.443$ $-0.321$ $0.000$ $-31.8$
2–3 vs 4–5 $-0.420$ $0.034$ $-0.486$ $-0.354$ $0.000$ $-34.3$
3–4 vs 5–6 $-0.727$ $0.127$ $-0.976$ $-0.478$ $0.000$ $-51.9$
4–5 vs 6–7 $-0.925$ $0.297$ $-1.507$ $-0.343$ $0.002$ $-60.3$
5–6 vs 7–8 $-0.877$ $0.399$ $-1.659$ $-0.095$ $0.028$ $-58.1$
6–7 vs 8–9 $-1.992$ $2.222$ $-6.346$ $2.363$ $0.370$ $-85.9$
7–8 vs 9–10 $-2.067$ $2.510$ $-6.987$ $2.854$ $0.401$ $-87.9$
Interpretation
The coefficient $\beta$ is interpreted as the semi-elasticity [^2] For a detailed discussion on accurately interpreting log-transformed variables in regression models, see (Huntington-Klein , 2023). of adjusted rank with respect to drafting disadvantage. Specifically, for binary $\text{Treat}_{it}$, the exponentiated coefficient gives the proportional change in adjusted rank:
%$\%\ \text{Change in Adjusted Rank}\approx\left(e^{\beta}-1\right)\times 100$
A negative $\beta$ thus implies that being in the latter drafting position group leads to a statistically significant improvement in performance (i.e., a lower adjusted rank). Early comparisons (e.g., 1–2 vs 3–4) yield large and precisely estimated effects, consistent with the hypothesis of a bandwagon drafting advantage. These effects attenuate and lose significance in later position comparisons, suggesting a capped or diminishing marginal benefit of drafting.
Pooled Bandwagon Effects (Lower Rank Pool = Treatment on Rank Outcome)
1–2 vs 3–42–3 vs 4–53–4 vs 5–64–5 vs 6–75–6 vs 7–81.21.2-1.2- 1.211-1- 10.80.8-0.8- 0.80.60.6-0.6- 0.60.40.4-0.4- 0.40.20.2-0.2- 0.200Drafting Position ComparisonEstimate (log pts)
Figure 3: Causal point estimates for the effect of latter drafter position on overall better rank in log pts. (significant comparisons only).

### v2

@@
 Focusing on open-water swim drafting—where athletes reduce drag most effectively by swimming directly behind peers—we estimate its performance effects through a structural contest framework with endogenous, deterministic effort and drafting position.
-Leveraging exogenous variation from COVID-19 drafting bans in Austrian triathlons, we apply a panel leave-one-out (LOO/LOTO) peer ability instrumental variables (IV) strategy to isolate the causal non-linear effect of drafting.
+Leveraging exogenous variation from COVID-19 drafting bans in Austrian triathlons, we apply a panel leave-one-out (LOO/LOTO) group ability instrumental variables (IV) strategy to isolate the causal non-linear effect of drafting.
 Results from restricted sample analysis and pooled estimated bandwagon IV effects show substantial and nonlinear gains: in small (group size $<$ 10) drafting swim groups/clusters, each deeper position improves finishing rank on average by over 30%, with rapidly diminishing returns in larger groups.
@@
 This paper investigates the performance implications of open water swim drafting using rich panel data from triathlon races by modeling drafting in the first segment as a two-stage deterministic game.
-In the first stage, athletes exert some deterministic swimming effort; in the second stage, they position themselves optimally within a group, potentially benefiting heterogeneously from reduced drag behind other swimmers due to the inversly decaying nature of hydrodynamic benefits and rare overtakings (resorting) during the swim segment (Reichel , 2025, 2025).
+In the first stage, athletes exert some deterministic swimming effort; in the second stage, they position themselves optimally within a group, potentially benefiting heterogeneously from reduced drag behind other swimmers due to the inversly decaying nature of hydrodynamic benefits and rare overtakings (resorting) during the swim segment (Reichel (2025)).
 Empirical estimates of drafting effects in the swim segment are still limited by measurement challenges.
 Group formations are likely endogenous and cannot be observed directly, and drafting exposure may correlate with unobserved ability (sorting/selection bias).
-Laboratory and biomechanical studies (e.g., Chatard , 1990, 2003; Delextrat , 2003; bolon2020drafting) focus on energy expenditure (metabolic) and performance gains from drafting, but large-scale causal estimates from race data have only been documented recently (in Reichel , 2025).
-In our theoretical framework, we extend contest theory and incorporate models of strategic externalities (Tullock , 1980; Moldovanu , 2001; Konrad , 2009; Che , 2003) to develop a structural model in which athletes first select a deterministic level of swimming effort, followed by an endogenous choice of drafting position.
+Laboratory and biomechanical studies (e.g., Chatard (1990, 2003); Delextrat (2003); Bolon (2020)) focus on energy expenditure (metabolic) and performance gains from drafting, but large-scale causal estimates from race data have only been documented recently (in Reichel (2025)).
+In our theoretical framework, we extend contest theory and incorporate models of strategic externalities (Tullock (1980); Moldovanu (2001); Konrad (2009); Che (2003)) to develop a structural model in which athletes first select a deterministic level of swimming effort, followed by an endogenous choice of drafting position.
 Drafting benefits are captured through a non-linear function that exhibits diminishing marginal returns beyond a certain positional threshold.
@@
 Earlier trials with complete linkage and wider thresholds (10–15 seconds) produced overly broad or unstable groupings.
-(Reichel , 2025)
+(Reichel (2025))
 Within each cluster, the fastest athlete is designated as the leader, and all others are assigned ordinal positions (e.g., second drafter, third drafter) based on swim-out time segment order.
@@
     A drafting position $d_{i}\in\mathbb{Z}_{\geq 1}$ within their drafting group in the second stage.
-This structure reflects a dynamic contest framework with endogenous positioning and group-related externalities (see Tullock , 1980; Konrad , 2009; Moldovanu , 2001; Tadelis , 2013; Vives , 2005).
+This structure reflects a dynamic contest framework with endogenous positioning and group-related externalities (see Tullock (1980); Konrad (2009); Moldovanu (2001); Tadelis (2013); Vives (2005)).
 3.1.1 Effort Cost Function
-For simplicity we assume a linear cost function of effort:
+Assumption 1.
+For the sake of simplicity we assume a linear cost function of effort:
 $C(e_{i}^{1})=c\cdot e_{i}^{1},\quad c>0$
@@
   • $\lambda>0$: rate of benefit growth.
-Proof:
+Lemma 1.
 For $d_{i}>3$, it can be shown that:
@@
 $\frac{dDIS_{i}}{dd_{i}}<0$
-Proof:
+Lemma 2.
 The exponential function is always positive, and $\gamma,\lambda,(1-\alpha)>0$, so the derivative is negative. Hence, disutility decreases with deeper position — but not indefinitely.
@@
 $\frac{d^{2}DIS_{i}}{dd_{i}^{2}}>0$
-Proof:
+Lemma 3.
 As the second derivative is positive, disutility is strictly convex in $d_{i}$, implying a unique minimum (interior optimum).
-3.2.3 Proposition (Interior Optimal Drafting Position):
-Proposition 1: There exists a unique $d_{i}^{*}>3$ such that disutility is minimized, i.e., drafting benefit is maximized net of performance shortfall risk.
-Proof: - For $d_{i}\leq 3$, $B(d_{i})=0\Rightarrow$ disutility is flat.
+Proposition 1 (Interior Optimal Drafting Position).
+There exists a unique $d_{i}^{*}>3$ such that disutility is minimized, i.e., drafting benefit is maximized net of performance shortfall risk.
+Proof.
+  • For $d_{i}\leq 3$, $B(d_{i})=0\Rightarrow$ disutility is flat.
   • For $d_{i}>3$, $DIS_{i}$ is strictly convex.
   • Since $DIS_{i}\to\alpha c\bar{e}_{i}^{1}-(1-\alpha)\gamma$ as $d_{i}\to\infty$, the marginal gains vanish.
-- Hence, the minimum is interior and finite.
+Hence, the minimum is interior and finite. ∎
 3.2.3 Empirical Mapping
@@
 $B(d_{i})=\gamma\left(1-e^{-\lambda(d_{i}-3)}\right)\cdot\mathbf{1}\{d_{i}>3\}$
-or to estimate effects. We can address address endogeneity by instrumenting $d_{i}$ using a leave-one-out peer ability IV.
+or to estimate effects. We can address address endogeneity by instrumenting $d_{i}$ using a leave-one-out group ability IV.
 3.3 Why the Optimal Position Is Not the Last
@@
 $\frac{dB}{dd_{i}}=\gamma\lambda e^{-\lambda(d_{i}-3)}\to 0$
-Conclusion: Athletes face trade-offs — drafting increases with depth, but so do tactical disadvantages.
+Corollary 1.
+Athletes face trade-offs — drafting increases with depth, but so do tactical disadvantages.
 Therefore, the optimal position is interior — balancing maximum benefit with control.
-This matches observed behavior in collective movement and contests with crowding externalities (Hillman , 1989; Krueger , 1974).
+This matches observed behavior in collective movement and contests with crowding externalities Hillman (1989); Krueger (1974).
 4 Identification Strategy and Empirical Framework
@@
 Because drafting positions reflect strategic sorting and group formation dynamics, naive OLS is likely biased.
-We thus use a two-stage least squares (2SLS) framework with a leave-one-out peer swim-time instrument to estimate a local average treatment effect (LATE) of drafting benefits (functional form from the theoretic model) or drafting positions on performance.
+We thus use a two-stage least squares (2SLS) framework with a leave-one-out group swim-time instrument to estimate a local average treatment effect (LATE) of drafting benefits (functional form from the theoretic model) or drafting positions on performance.
 4.2 Structural Potential Outcomes Framework
@@
 $\text{Cov}(D_{ieg},\varepsilon_{ie})\neq 0$
-We address this using a constructed leave-one-out peer swim-time instrument $Z_{ie}$ to generate the relevant and exogenous variation in swim group drafting structure.
+We address this using a constructed leave-one-out group swim-time instrument $Z_{ie}$ to generate the relevant and exogenous variation in swim group drafting structure.
 4.3.1 Leave-One-Out (LOO/LOTO) Instrumental Variable (IV)
-We, define the instrument as:
+Definition 1 (LOO/LOTO Instrument).
+We define the instrument as:
 $Z_{ie}=\frac{1}{|\mathcal{G}_{ie}^{-i}|}\sum_{j\in\mathcal{G}_{ie}^{-i}}S_{je}$
-This captures the swim ability of peers in athlete $i$’s drafting within the current group, excluding $i$.This leave-one-out (LOO/LOTO) instrument captures the average peer ability while avoiding mechanical reflection bias.
+This captures the swim ability of groups in athlete $i$’s drafting within the current group, excluding $i$.
+This leave-one-out (LOO/LOTO) instrument captures the average group ability while avoiding mechanical reflection bias.
 By excluding athlete $i$, we remove the mechanical correlation between the instrument and the outcome.
-This strategy aligns with methods used in social interaction and peer effect models to address endogeneity, e.g., Bramoulle (2009) and Angrist (2014).
-Related applications include De Loecker (2011), who employ LOO instruments in production function estimation, and Guryan (2008), who instrument for peer characteristics using group averages excluding the individual.
+This strategy aligns with methods used in social interaction and group effect models to address endogeneity, e.g., Bramoulle (2009); Angrist (2014).
+Related applications include De Loecker (2011), who employ LOO instruments in production function estimation, and Guryan (2008), who instrument for group characteristics using group averages excluding the individual.
 4.3.2 Drafting Benefit Transformation from Theoretic Model
@@
 This transformation captures the increasing-but-saturating drafting benefits posited by the theoretical model.
-4.4 Empirical Specification: Two-Stage Least Squares (2SLS) Peer-IV Estimation
-To identify the causal effect of within-group drafting benefit on race performance, we estimate a two-stage least squares (2SLS) model using a leave-one-out peer ability instrument.
+4.4 Empirical Specification: Two-Stage Least Squares (2SLS) group-IV Estimation
+To identify the causal effect of within-group drafting benefit on race performance, we estimate a two-stage least squares (2SLS) model using a leave-one-out group ability instrument.
 This setup accounts for endogeneity in observed drafting positions, which are determined by both athlete ability and strategic group formation.
@@
 First Stage: Instrumenting Drafting Position
-We begin by estimating the drafting position as a function of peer swim ability and controls:
+We begin by estimating the drafting position as a function of group swim ability and controls:
 $D_{ieg}=\pi_{0}+\pi_{1}Z_{ie}+\pi_{2}S_{ie}+\bm{X}_{ie}^{\prime}\bm{\pi}_{3}+\lambda_{e}+\alpha_{i}+\alpha_{g}+u_{ieg}$
@@
   • $D_{ieg}$: observed drafting position of athlete $i$ in event $e$, group $g$;
-  • $Z_{ie}$: leave-one-out peer average swim time (excluding athlete $i$);
+  • $Z_{ie}$: leave-one-out group average swim time (excluding athlete $i$);
   • $S_{ie}$: own swim time of athlete $i$, capturing own ability;
@@
   • $u_{ieg}$: idiosyncratic first-stage error term.
-The instrument $Z_{ie}$ introduces exogenous variation by exploiting group composition.
-Since athletes do not control who else is in their drafting cluster, this measure captures plausibly exogenous shifts in strategic positioning pressure.
 Second Stage: Estimating the Effect of Drafting Benefit on Performance
-Using the fitted drafting benefit $\widehat{B}_{ieg}=f(\widehat{D}_{ieg})$ to validate the theoretical model or using $\widehat{D}_{ieg})$ directly, we estimate:
+Using the fitted drafting benefit $\widehat{B}_{ieg}=f(\widehat{D}_{ieg})$ to validate the theoretical model or using $\widehat{D}_{ieg}$ directly, we estimate:
 $\log(rank_{ie}+1)=\beta_{1}\widehat{B}_{ieg}+\beta_{2}\cdot leader_{ieg}+\bm{X}_{ie}^{\prime}\bm{\beta}_{3}+\lambda_{e}+\alpha_{i}+\alpha_{g}+\varepsilon_{ieg}$
@@
   • $\varepsilon_{ieg}$: second-stage error term.
-Identification Strategy and Exogeneity
-This 2SLS framework identifies the local average treatment effect (LATE) of drafting benefit for compliers—athletes whose position is affected by peer group composition. Identifying Assumptions are:
+4.4.1 Identification Strategy and Exogeneity
+This 2SLS framework identifies the local average treatment effect (LATE) of drafting benefit for compliers—athletes whose position is affected by group group composition.
+Assumption 2 (Identification Assumptions).
   • Instrument relevance: $\text{Cov}(Z_{ie},D_{ieg})\neq 0$;
-  • Instrument exogeneity: $\mathbb{E}[\varepsilon_{ieg}\mid Z_{ie},S_{ie},\bm{X}_{ie},\lambda_{e},\alpha_{i},\alpha_{g}]=0$. – $Z_{ie}$;
-  • Monotonicity: An increase in peer ability (faster peers) weakly shifts athletes to deeper drafting positions.
+  • Instrument exogeneity: $\mathbb{E}[\varepsilon_{ieg}\mid Z_{ie},S_{ie},\bm{X}_{ie},\lambda_{e},\alpha_{i},\alpha_{g}]=0$;
+  • Monotonicity: An increase in group ability (faster groups) weakly shifts athletes to deeper drafting positions.
 This empirical structure provides consistent estimates of the causal effect of drafting benefit on race outcomes while accounting for sorting, selection, and unobserved heterogeneity through fixed effects and instrument design.
-where: - $\widehat{B}_{ieg}$: instrumented drafting benefit; - $leader_{ieg}$: indicator for leading position ($D=1$); - $\lambda_{e},\alpha_{i},\alpha_{g}$: event, athlete, and group fixed effects.
 4.4.2 Structural Validation and Interpretation
-The coefficient $\beta_{1}$ reflects the average marginal return to drafting benefit.
-The increasing and concave structure implies that: - Greatest positive net gains occur when transitioning from front positions backwards; Diminishing returns occur deeper into the pack; The optimal position is interior, as predicted by the theoretical model.
-Empirical evidence of this shape supports the game’s interior equilibrium logic and nonlinear (dis-)utility (drafting befenefit structure) topology.
+The coefficient $\beta_{1}$ reflects the average marginal return to drafting benefit. The increasing and concave structure implies that:
+  • Greatest positive net gains occur when transitioning from front positions backwards;
+  • Diminishing returns occur deeper into the pack;
+  • The optimal position is interior, as predicted by the theoretical model.
+Empirical evidence of this shape supports the game’s interior equilibrium logic and nonlinear (dis-)utility (drafting benefit structure) topology.
 4.5 Potential Extension: Projected Panel IV
-For robustness and within-athlete variation, we define a projected instrument using historical group characteristics from the largest group containing athlete ’i’ f.e.:
+For robustness and within-athlete variation, we define a projected instrument using historical group characteristics from the largest group containing athlete $i$, e.g.:
 $Z_{ie^{\prime}}^{\text{histProjected}}=\mathbb{E}_{e}[Z_{ie}\mid i\in g_{ie},e\neq e^{\prime}]$
@@
 The instrumental variables (IV) estimates consistently show that deeper drafting positions improve performance, with stronger estimated effects in smaller groups.
-In Table 3, the coefficient on drafting position is $-0.011$ ($p<0.1$) in the full sample (without restricting the outcome variable), implying that moving one position deeper in the drafting line improves expected rank by approximately $1.1\%$:
+Proposition 2.
+In Table 3, the coefficient on drafting position is $-0.011$ ($p<0.1$) in the full sample (without restricting the outcome variable), implying that moving one position deeper in the drafting line improves expected rank by approximately 1.1%:
 %$\%\Delta r_{i}\approx(e^{-0.011}-1)\times 100\approx-1.1\%$
 This effect increases substantially in smaller groups.
-In drafting clusters with fewer than 10 athletes, the coefficient becomes $-0.972$ ($p<0.01$), translating to a $62.1\%$ improvement in finishing rank per position:
+In drafting clusters with fewer than 10 athletes, the coefficient becomes $-0.972$ ($p<0.01$), translating to a 62.1% improvement in finishing rank per position:
 %$\%\Delta r_{i}\approx(e^{-0.972}-1)\times 100\approx-62.1\%$
-In very small clusters (group size $<5$), the coefficient is $-2.378$ ($p<0.1$), corresponding to an $90.7\%$ improvement:
+In very small clusters (group size $<5$), the coefficient is $-2.378$ ($p<0.1$), corresponding to a 90.7% improvement:
 %$\%\Delta r_{i}\approx(e^{-2.378}-1)\times 100\approx-90.7\%$
@@
 $B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ 1-e^{-\lambda(d_{i}-3)},&\text{if }d_{i}>3\end{cases}$
-with some uncalibrated $\lambda=0.5$ for the simple sake of simplicity and brevity.
-This specification reflects increasing but diminishing returns to deeper drafting positions.
+with some uncalibrated $\lambda=0.5$ for the simple sake of simplicity and brevity. This specification reflects increasing but diminishing returns to deeper drafting positions.
+Observation 1.
 Across all models, the coefficient on $B(d_{i})$ is significantly negative, consistent with improved performance from stronger drafting effects.
@@
 Given the IV framework controls for endogenous group sorting, the leader coefficient likely captures small/negligible non-effort-based advantages such as tactical control, friction or reduced group congestion—complementing prior evidence on physiological pacing costs (Menting , 2022).
+Proposition 3.
 All models reject exogeneity via the Wu–Hausman test ($p<2.2\times 10^{-16}$), validating the use of instrumental variables.
@@
 This paper contributes comprehensive analysis on Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-linear Externalities and draws on empirical causal evidence using rich panel data from drafting in endurance sports competitions, predicting effects on final stage rank outcomes focusing on competitive athletes.
-We develop and estimate log-linear models with endogenous drafting position, using a leave-one-out peer swim ability as an instrument (Angrist , 2009; Manski , 1993; De Giorgi , 2020).
+We develop and estimate log-linear models with endogenous drafting position, using a leave-one-out group swim ability as an instrument (Angrist , 2009; Manski , 1993; degiorgi2020group).
 The results validate the strategic and nonlinear nature of drafting benefits in multi-stage competitions such as triathlon.
-Our key findings are:
   1. 1.
@@
 Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-Linear Externalities: Evidence from Triathlon using a Leave-One-Out IV Strategy
-A.1 Leave-one-out Peer-IV 2SLS Regressions with Capped/Limited Outcome Variable Rank $<$ 250
+A.1 Leave-one-out Group-IV 2SLS Regressions with Capped/Limited Outcome Variable Rank $<$ 250
 Table 2: Instrumental Variables Estimates of Drafting Effects on Performance
@@
 $R^{2}$ 0.641 -0.166 -0.674 Within $R^{2}$0.905 -5.925 -8.453 First Stage F-stat 1,115.6 51.4 12.7 First Stage $p$-value $<2.2\times 10^{-16}$ $8.23\times 10^{-13}$ $3.64\times 10^{-4}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the centered finishing rank of athlete $i$ within event.
-The endogenous regressor is the athlete’s drafter position $p_{i}$, instrumented using the cluster-level peer mean swim performance (leave-one-out).
+The endogenous regressor is the athlete’s drafter position $p_{i}$, instrumented using the cluster-level Group mean swim performance (leave-one-out).
 No additional covariates are included.
@@
 Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
-A.2 Robustness: Leave-one-out Peer-IV 2SLS Regressions under different sample sizes and model specifications
+A.2 Robustness: Leave-one-out Group-IV 2SLS Regressions under different sample sizes and model specifications
 Table 4: The Causal Effects of Drafting Position on Log Race Rank (2SLS Estimates)
@@
 Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
-A.3 Robustness: Leave-one-out Peer-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function
+A.3 Robustness: Leave-one-out Group-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function
 Table 5: Instrumental Variables Estimates of Theory-Based Drafting Effects on Performance
@@
 The endogenous variable is a theory-based drafting benefit: $B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ \gamma\left(1-e^{-\lambda(d_{i}-3)}\right),&\text{if }d_{i}>3\end{cases}$ with $\gamma=1$ and $\lambda=0.5$.
-The instrument is the cluster-level peer mean swim time (leave-one-out).
+The instrument is the cluster-level Group mean swim time (leave-one-out).
 Fixed effects for athlete, cluster, and event are included.
@@
 The endogenous regressor is the theory-based drafting benefit: $B(d_{i})=1-e^{-\lambda(d_{i}-3)}$ for $d_{i}>3$, and 0 otherwise, with $\lambda=0.5$.
-Instrument: leave-one-out peer swim time.
+Instrument: leave-one-out Group swim time.
 Fixed effects for athlete, event, and cluster are included.
@@
 Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$
-Appendix B Data and Descriptive Statistics as in (Reichel , 2025)
+Appendix B Data and Descriptive Statistics (as in (Reichel , 2025))
 Variable Min 1st Quartile Median Mean 3rd Quartile Max Total 165 4965 8232 11789 16809 61282 Swim-Out Times 4 825 1431 1625 2097 29700 Race Rank 0 28 69 156.9 151 2623 Male 0.0000 1.0000 1.0000 0.7898 1.0000 1.0000 Year 1922 1969 1977 1977 1985 2009 Covid 0.0000 0.0000 0.0000 0.1097 0.0000 1.0000 Post Covid 0.00000 0.00000 0.00000 0.08563 0.00000 1.00000 Week Running 0.0 159.0 279.0 317.6 475.0 751.0 Age 8.00 31.00 38.00 38.56 46.00 99.00 Age Squared 64 961 1444 1600 2116 9801 Event Year 2010 2013 2015 2016 2019 2024 Cluster (Swim Group) 1.00 4.00 10.00 16.95 23.00 224.00 Leader 0.0000 0.0000 0.0000 0.3134 1.0000 1.0000 Drafter 0.0000 0.0000 1.0000 0.6866 1.0000 1.0000 Drafter Position 1.00 1.00 3.00 17.15 11.00 937.00 First Drafter 0.0000 0.0000 0.0000 0.2961 1.0000 1.0000 Second Drafter 0.0000 0.0000 0.0000 0.1342 0.0000 1.0000 Third Drafter 0.00000 0.00000 0.00000 0.08329 0.00000 1.00000 Fourth Drafter 0.00000 0.00000 0.00000 0.05846 0.00000 1.00000 Fifth Drafter 0.000 0.000 0.000 0.044 0.000 1.000 Last Drafter 0.0000 0.0000 0.0000 0.2961 1.0000 1.0000 Notes: This table presents summary statistics for the key variables used in the analysis.
@@
   • $\varepsilon_{it}$ captures unobserved performance shocks.
-Because $\text{Treat}_{it}$ may be endogenous—potentially correlated with $\varepsilon_{it}$—we instrument for it using the exogenous variation in peer group swim ability. The first-stage equation is:
-$\text{Treat}_{it}=\pi\cdot\text{PeerSwimMean}_{it}+\eta_{i}+\theta_{t}+u_{it}$ (2)
+Because $\text{Treat}_{it}$ may be endogenous—potentially correlated with $\varepsilon_{it}$—we instrument for it using the exogenous variation in Group group swim ability.
+The first-stage equation is:
+$\text{Treat}_{it}=\pi\cdot\text{GroupSwimMean}_{it}+\eta_{i}+\theta_{t}+u_{it}$ (2)
 where:
-  • $\text{PeerSwimMean}_{it}$ is the leave-one-out average swim score of athlete $i$’s drafting cluster in event $t$.
+  • $\text{GroupSwimMean}_{it}$ is the leave-one-out average swim score of athlete $i$’s drafting cluster in event $t$.
   • $\eta_{i}$, $\theta_{t}$ are the same fixed effects, and $u_{it}$ is the first-stage error term.
-Under the standard IV assumptions—instrument relevance ($\pi\neq 0$), instrument exogeneity ($\mathbb{E}[\text{PeerSwimMean}_{it}\cdot\varepsilon_{it}]=0$), and monotonicity—the parameter $\beta$ captures the local average treatment effect (IV-LATE) of being a drafter in latter pooled drafter position band on log-transformed, adjusted rank.
+Under the standard IV assumptions—instrument relevance ($\pi\neq 0$), instrument exogeneity ($\mathbb{E}[\text{GroupSwimMean}_{it}\cdot\varepsilon_{it}]=0$), and monotonicity—the parameter $\beta$ captures the local average treatment effect (IV-LATE) of being a drafter in latter pooled drafter position band on log-transformed, adjusted rank.
 Results